OpenStudy (anonymous):

What is the area of a square with perimeter 4x – 12?

OpenStudy (anonymous):

first you need to get the length of a side of the square... since the perimeter i given, divide that expression by 4.. what you have left is the length of a side of the square. to get the area, square that expression.

OpenStudy (anonymous):

it will give you 3

OpenStudy (anonymous):

divide the perimeter by 4: \(\large \frac{4x-12}{4}= \) ????

OpenStudy (anonymous):

@dpalnc how can I do that. Please help me.. do I cancel out anything or multiply -12*4?

OpenStudy (anonymous):

\[\huge \frac{4x-12}{4}=\frac{4(x-3)}{4}=\frac{\cancel4^1(x-3)}{\cancel4^1}=x-3 \]

OpenStudy (lgbasallote):

it's dpaInc not dpanlc :S

OpenStudy (lgbasallote):

like this @dpaInc

OpenStudy (anonymous):

lol

OpenStudy (lgbasallote):

no such dpalnc exists in openstudy..just a tip ;)

OpenStudy (anonymous):

so x-3 represent the length of a side of the square... to get the area, square that binomial...

OpenStudy (anonymous):

@dpaInc okay.. let me do that

OpenStudy (anonymous):

x-6x+9???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw...:)

OpenStudy (callisto):

Very little correction over there.. Area is \(x^2-6x+9\)

OpenStudy (anonymous):

sorry..., @Callisto ,...:)

OpenStudy (callisto):

Nope.. no worries!!~

OpenStudy (anonymous):

i need 4 letters to tag u now....

OpenStudy (anonymous):

it used to be only 2...

OpenStudy (anonymous):

lol. I see what happened there ;)

OpenStudy (anonymous):

sorry folks.... i'll be careful.....